OpenStudy (anonymous):

need help solving this equation (x-4)(2x+5)=0

OpenStudy (anonymous):

you want to obtain the values of x?

OpenStudy (anonymous):

The thing is either of the two expressions x-4=0 or 2x+5=0 x=4 or 2x=-5, i.e. x=4, or x=-5/2

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

(x-4)(2x+5)=0 Multiply each term of inside the first parenthesis to each term in the second parenthesis and add: 2x^2 + 5x - 8x -20 = 0 2x^2 + 13x -20 = 0 thiis is a quadratic equation you can use the formula: x = [-b +- (b^2 - 4ac)^1/2] / 2a a = 2, b = 13, c = 20 Remember, you will get two solutions x1 and x2 x1 = [-2 + (169 - 4*2*20)^1/2]/4 = [-2 + (169 - 160)^1/2]/4 = (-2 + 9^1/2)/4 = (-2 + 3)/4 = 1/4 x2 = [-2 - (169 - 4*2*20)^1/2]/4 = [-2 - (169 - 160)^1/2]/4 = (-2 -3)/4 = -5/4 Sorry 2x^2 + 5x - 8x -20 = 0 is 2x^2 - 3x -20 = 0 the quadratic equation is: 2x^2 - 3x -20 = 0 a = 2 b= -3 c= -20

OpenStudy (anonymous):